# Adversarial Training on Purification (AToP): Advancing Both Robustness and Generalization

### v3

@@ Abstract
The deep neural networks are known to be vulnerable to well-designed adversarial attacks. The most successful defense technique based on adversarial training (AT) can achieve optimal robustness against particular attacks but cannot generalize well to unseen attacks. Another effective defense technique based on adversarial purification (AP) can enhance generalization but cannot achieve optimal robustness. Meanwhile, both methods share one common limitation on the degraded standard accuracy. To mitigate these issues, we propose a novel pipeline to acquire the robust purifier model, named Adversarial Training on Purification (AToP), which comprises two components: perturbation destruction by random transforms (RT) and purifier model fine-tuned (FT) by adversarial loss. RT is essential to avoid overlearning to known attacks, resulting in the robustness generalization to unseen attacks, and FT is essential for the improvement of robustness. To evaluate our method in an efficient and scalable way, we conduct extensive experiments on CIFAR-10, CIFAR-100, and ImageNette to demonstrate that our method achieves optimal robustness and exhibits generalization ability against unseen attacks.
@@ 1 Introduction
[Uncaptioned image] Figure 1: Illustration of adversarial training on purification (AToP).
@@ 4.1 Experimental setup
Table 2: Standard accuracy and robust accuracy against AutoAttack $l_{\infty}$ threat ($\epsilon=8/255$) on CIFAR-10 with WideResNet classifier model. Defense method Extra Standard Robust data Acc. Acc.  WideResNet-28-10 Zhang et al. (2020) 85.36 59.96 Gowal et al. (2020) 89.48 62.70 Rebuffi et al. (2021) $\times$ 87.33 61.72 Gowal et al. (2021) $\times$ 87.50 65.24 Nie et al. (2022) $\times$ 89.02 70.64  Ours $\times$ 90.62 72.85  WideResNet-70-16 Gowal et al. (2020) 91.10 65.87 Rebuffi et al. (2021) 92.23 66.56 Rebuffi et al. (2021) $\times$ 88.54 64.46 Gowal et al. (2021) $\times$ 88.74 66.60 Nie et al. (2022) $\times$ 90.07 71.29  Ours $\times$ 91.99 76.37 Table 5: Robust accuracy against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) on CIFAR-10 and CIFAR-100,AutoAttack $l_{2}$ ($\epsilon=0.5$) on CIFAR-10 with WideResNet-28-10 classifier model. Defense method CIFAR CIFAR CIFAR 10, $l_{\infty}$ 10, $l_{2}$ 100, $l_{\infty}$ Wu et al. (2023) 44.73 45.12 38.48 Wu et al. (2023) 59.18 65.43 40.23  +AToP (Ours) Ughini et al. (2022) 59.57 63.69 38.89 Ughini et al. (2022) 72.85 80.47 43.55  +AToP (Ours)
Table 3: Standard accuracy and robust accuracy against AutoAttack $l_{2}$ threat ($\epsilon=0.5$) on CIFAR-10 with WideResNet classifier model. Defense method Extra Standard Robust data Acc. Acc.  WideResNet-28-10 Augustin et al. (2020) 92.23 77.93 Ding et al. (2019) $\times$ 88.02 67.77 Rebuffi et al. (2021) $\times$ 91.79 78.32 Nie et al. (2022) $\times$ 91.03 78.58  Ours $\times$ 90.62 80.47  WideResNet-70-16 Rebuffi et al. (2021) 95.74 81.44 Gowal et al. (2020) $\times$ 90.90 74.03 Nie et al. (2022) $\times$ 92.68 80.60  Ours $\times$ 91.99 81.35 Table 4: Standard accuracy and robust accuracy against AutoAttack $l_{\infty}$ threat ($\epsilon=8/255$) on CIFAR-100 with WideResNet classifier model. Defense method Extra Standard Robust data Acc. Acc.  WideResNet-28-10 Hendrycks et al. (2019) 59.23 28.42 Pang et al. (2022) $\times$ 63.66 31.08 Cui et al. (2023) $\times$ 73.85 39.18  Ours $\times$ 74.61 43.55  WideResNet-70-16 Bai et al. (2023) 85.21 38.72 Wang et al. (2023) $\times$ 75.22 42.67  Ours $\times$ 76.95 44.44
@@ 4.2 Comparison with state-of-the-art methods
Table 5 shows the performance of the two generator-based purifier models (Ughini et al., 2022; Wu et al., 2023) against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) and $l_{2}$ ($\epsilon=0.5$) threat models with WideResNet-28-10 classifier model. Specifically, our method improves the robust accuracy by 14.45% and 20.31% on AE-based purifier model, 13.28% and 16.87% on GAN-based purifier model, respectively. The results demonstrate that our method can significantly improve the robustness of the purifier model.
Result analysis on CIFAR-100: Table 4 shows the performance of the defense methods against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) threat model on CIFAR-100 with WideResNet-28-10 and WideResNet-70-16 classifier models. Our method outperforms all other methods without extra data in terms of both standard accuracy and robust accuracy. Specifically, compared to the second-best method, our method improves the robust accuracy by 4.37% on WideResNet-28-10 and by 1.77% on WideResNet-70-16, respectively. Table 5 shows the performance of the models against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) and the observations are basically consistent with CIFAR-10, demonstrating the effectiveness of our method in enhancing robust classification across different datasets.
@@ 4.4 Ablation studies
Table 8: Standard accuracy and robust accuracy against PGD+EOT $l_{2}$ threat ($\epsilon=1.0$) on CIFAR-10 with ResNet-18 classifier model. AToP Standard PGD-10 PGD-20 $\rm{RT_{1}}$$\times$ 93.36 20.31 15.04 93.36 27.34 22.46 $\rm{RT_{2}}$$\times$ 84.18 67.19 63.09 90.04 69.92 68.75 $\rm{RT_{3}}$$\times$ 75.98 28.52 26.17 80.02 46.48 41.60
@@
Limitations: Given that our method involves the fine-tuning of the generative model, the complexity of the generative model has a direct impact on the computational cost of AToP. Initially, we explore fine-tuning the diffusion model to improve its robustness. However, as purified images need to be fed into the classifier, generating 1000 images and fine-tuning the diffusion model requires 144 minutes. In contrast, with the same setting, the GAN-based purifier model only requires 62 seconds.
@@ 5 Conclusion
In the paper, we propose a novel defense technique called AToP that utilizes adversarial training to acquire the robust purifier model. To demonstrate the robustness of our method, we conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette. In defense of various attacks under multiple transforms, classifier and purifier architectures, our method consistently achieves state-of-the-art results and exhibits generalization ability against unseen attacks. Ablation studies further highlight that our method can significantly improve the performance of the purifier model on robust classification. Despite the large improvements achieved, our method has a major limitation: AToP requires training on the purifier model, and as the complexity of the purifier model increases, so does the training cost. However, exploring this direction presents an intriguing research opportunity for developing a more efficient defense method by combining AT and AP.

### v4

@@ Abstract
The deep neural networks are known to be vulnerable to well-designed adversarial attacks. The most successful defense technique based on adversarial training (AT) can achieve optimal robustness against particular attacks but cannot generalize well to unseen attacks. Another effective defense technique based on adversarial purification (AP) can enhance generalization but cannot achieve optimal robustness. Meanwhile, both methods share one common limitation on the degraded standard accuracy. To mitigate these issues, we propose a novel pipeline to acquire the robust purifier model, named Adversarial Training on Purification (AToP), which comprises two components: perturbation destruction by random transforms (RT) and purifier model fine-tuned (FT) by adversarial loss. RT is essential to avoid overlearning to known attacks, resulting in the robustness generalization to unseen attacks, and FT is essential for the improvement of robustness. To evaluate our method in an efficient and scalable way, we conduct extensive experiments on CIFAR-10, CIFAR-100, and ImageNette to demonstrate that our method achieves optimal robustness and exhibits generalization ability against unseen attacks. Our code is available at https://github.com/glin2022/atop.
@@ 1 Introduction
Figure 1: Illustration of adversarial training on purification (AToP).
Table 1: Robustness comparison of defenses with expectation (negative impacts are marked in red).
Defense Clean Known Unseen
method images attacks attacks
[…]
AP $\downarrow$ $\uparrow$ $\uparrow$
@@ 4.1 Experimental setup
Table 2: Standard accuracy and robust accuracy against AutoAttack $l_{\infty}$ threat ($\epsilon=8/255$) on CIFAR-10 with WideResNet classifier model.
Defense method Extra Standard Robust
data Acc. Acc.
WideResNet-28-10
Zhang et al. (2020) 85.36 59.96
Gowal et al. (2020) 89.48 62.70
Rebuffi et al. (2021) $\times$ 87.33 61.72
Gowal et al. (2021) $\times$ 87.50 65.24
Nie et al. (2022) $\times$ 89.02 70.64
Ours $\times$ 90.62 72.85
WideResNet-70-16
Gowal et al. (2020) 91.10 65.87
Rebuffi et al. (2021) 92.23 66.56
Rebuffi et al. (2021) $\times$ 88.54 64.46
Gowal et al. (2021) $\times$ 88.74 66.60
Nie et al. (2022) $\times$ 90.07 71.29
Ours $\times$ 91.99 76.37
Table 5: Robust accuracy against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) on CIFAR-10 and CIFAR-100,AutoAttack $l_{2}$ ($\epsilon=0.5$) on CIFAR-10 with WideResNet-28-10 classifier model.
Defense method CIFAR CIFAR CIFAR
10, $l_{\infty}$ 10, $l_{2}$ 100, $l_{\infty}$
Wu et al. (2023) 44.73 45.12 38.48
Wu et al. (2023) 59.18 65.43 40.23
+AToP (Ours)
Ughini et al. (2022) 59.57 63.69 38.89
Ughini et al. (2022) 72.85 80.47 43.55
+AToP (Ours)
Table 3: Standard accuracy and robust accuracy against AutoAttack $l_{2}$ threat ($\epsilon=0.5$) on CIFAR-10 with WideResNet classifier model.
Defense method Extra Standard Robust
data Acc. Acc.
WideResNet-28-10
Augustin et al. (2020) 92.23 77.93
Ding et al. (2019) $\times$ 88.02 67.77
Rebuffi et al. (2021) $\times$ 91.79 78.32
Nie et al. (2022) $\times$ 91.03 78.58
Ours $\times$ 90.62 80.47
WideResNet-70-16
Rebuffi et al. (2021) 95.74 81.44
Gowal et al. (2020) $\times$ 90.90 74.03
Nie et al. (2022) $\times$ 92.68 80.60
Ours $\times$ 91.99 81.35
Table 4: Standard accuracy and robust accuracy against AutoAttack $l_{\infty}$ threat ($\epsilon=8/255$) on CIFAR-100 with WideResNet classifier model.
Defense method Extra Standard Robust
data Acc. Acc.
WideResNet-28-10
Hendrycks et al. (2019) 59.23 28.42
Pang et al. (2022) $\times$ 63.66 31.08
Cui et al. (2023) $\times$ 73.85 39.18
Ours $\times$ 74.61 43.55
WideResNet-70-16
Bai et al. (2023) 85.21 38.72
Wang et al. (2023) $\times$ 75.22 42.67
Ours $\times$ 76.95 44.44
@@ 4.2 Comparison with state-of-the-art methods
Table 4.1 shows the performance of the two generator-based purifier models (Ughini et al., 2022; Wu et al., 2023) against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) and $l_{2}$ ($\epsilon=0.5$) threat models with WideResNet-28-10 classifier model. Specifically, our method improves the robust accuracy by 14.45% and 20.31% on AE-based purifier model, 13.28% and 16.87% on GAN-based purifier model, respectively. The results demonstrate that our method can significantly improve the robustness of the purifier model.
Result analysis on CIFAR-100: Table 4 shows the performance of the defense methods against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) threat model on CIFAR-100 with WideResNet-28-10 and WideResNet-70-16 classifier models. Our method outperforms all other methods without extra data in terms of both standard accuracy and robust accuracy. Specifically, compared to the second-best method, our method improves the robust accuracy by 4.37% on WideResNet-28-10 and by 1.77% on WideResNet-70-16, respectively. Table 4.1 shows the performance of the models against AutoAttack $l_{\infty}$ ($\epsilon=8/255$) and the observations are basically consistent with CIFAR-10, demonstrating the effectiveness of our method in enhancing robust classification across different datasets.
@@ 4.4 Ablation studies
Table 8: Standard accuracy and robust accuracy against PGD+EOT $l_{2}$ threat ($\epsilon=1.0$) on CIFAR-10 with ResNet-18 classifier model.
AToP Standard PGD-10 PGD-20
$\rm{RT_{1}}$$\times$ 93.36 20.31 15.04
93.36 27.34 22.46
$\rm{RT_{2}}$$\times$ 84.18 67.19 63.09
90.04 69.92 68.75
$\rm{RT_{3}}$$\times$ 75.98 28.52 26.17
80.02 46.48 41.60
@@
Limitations: Given that our method involves the fine-tuning of the generative model, the complexity of the generative model has a direct impact on the computational cost of AToP. Initially, we explore fine-tuning the diffusion model to improve its robustness. However, as purified images need to be fed into the classifier, generating 1000 images and fine-tuning the diffusion model requires 144 minutes. In contrast, with the same setting, the GAN-based purifier model only requires 62 seconds. To address this issue, we propose a new robust diffusion model for adversarial purification, more details in our another paper (Lin et al., 2024).
@@ 5 Conclusion
In the paper, we propose a novel defense technique called AToP that utilizes adversarial training to acquire the robust purifier model. To demonstrate the robustness of our method, we conduct extensive experiments on CIFAR-10, CIFAR-100 and ImageNette. In defense of various attacks under multiple transforms, classifier and purifier architectures, we find the pre-trained generator is not good enough for classification and robustness, and our method consistently achieves state-of-the-art results and exhibits generalization ability against unseen attacks. Despite the large improvements achieved, our method has a major limitation: AToP requires training on the purifier model, and as the complexity of the purifier model increases, so does the training cost. However, exploring this direction presents an intriguing research opportunity to develop a more efficient defense method by combining AT and AP.
Acknowledgments
Guang Lin was supported by the RIKEN Junior Research Associate Program. This work was supported in part by the JSPS KAKENHI Grant Numbers JP20H04249, JP23K28109, 24K03005, and RIKEN Incentive Research Project.